1.The electricity costs of a business decreased from 12,00 one year to 10,920 the next year what is the rate of decrease?


2.The total bill was $65.32 with the 15% gratuity already added what was the cost of the actual meal?

Found 2 solutions by awhalen, tsmith5651:
Answer by awhalen(4)

1.10290=12000(1-r).find the r. r=10290/12000= .91
1-.91= .09 to get a percentage multiply .09 by 100
9% decrease rate
2. 65.32*.15= 9.798 add 9.798 to the original cost.
65.32+9.798= 75.118 round to the nearest tenth $75.12


Answer by tsmith5651(7)

For problem #2.
are you sure this is what the math problem is asking for? My beginning algebra book asks it this way.Section 2.6 problem 58 If it is the same as below. I have given you what they are looking for.
If the total bill at a restaurant, including a 15% tip is $65.32, what was the cost of the meal alone?

Meal Cost = 100%m = m
Tip Cost = 15% = 0.15m
m+0.15m=65.32
1.15m=65.32
m=56.80
$original cost of meal $56.80
Check* 56.80 x 15.%=$65.32